what is the kinetic energy of a ball with a mass of 5 kg rolling at 10 m/s
Answer #1
Total kinetic energy depends upon whether the ball is solid or hollow. Total kinetic energy is the sum of translational kinetic energy (Etran),
(M/2)*V^2, and rotational kinetic energy (Erot). The rotational kinetic energy depends upon the moment of inertia.
Erot = (1/2)I*w^2
where w = V/R.

For a solid ball of uniform density,

I = (2/5)M*R^2, and
Erot = (1/2)(2/5)M*R^2*(V/R)^2
= (M/5) V^2

Etotal = Etrans + Erot = (7/10)M*V^2
answered by: Learner